OpenStudy (anonymous):

Solve the given equation. 12/v^2-16-24/v-4=3

OpenStudy (anonymous):

\[\frac{ 12 }{ v^2-16 }-\frac{24 }{ v-4 }=3\] \[\frac{ 12-24(v+4) }{v^2-16 }=3\] \[\left\{ v^2-16=\left( v+4 \right)\left( v-4 \right) \right\}\] \[12-24v-96=3(v^2-16)\] \[3v^2+24v-48-12+96=0\] \[3v^2+24v+36=0,v^2+8v+12=0,v^2+2v+6v+12=0\] make factors and find the value of v.